OpenStudy (anonymous):

HELP PLEASE!<3 ANSWER WILL GET MEDAL! Part 1: When writing linear equations, how do you determine which form of a line to use? Part 2: Choose 1 set of points from the choices below. Then, solve the problem and post your solution, showing your steps. Write an equation in point-slope form for the line that passes through one of the following pairs of points (you may choose the pair you want to work with). Then, use the same set of points to write the equation in standard form and again in slope-intercept form. Point pairs ◦(5, 1), (–3, 4) ◦(0, –2), (3, 2) ◦(–2, –1), (1, 2)

OpenStudy (johnweldon1993):

Hmm well it depends on what information we have available... if we have a point...and a slope...makes sense we would use the point slope form if we have just the slope and say a y-intercept...we would use slope intercept form Standard form would be good when we want to find out the intercepts right away...

OpenStudy (johnweldon1993):

Now lets choose a point...pick one :)

OpenStudy (anonymous):

◦(5, 1), (–3, 4)

OpenStudy (johnweldon1993):

Alright so we have a point first step....find the slope \[\large m=\frac{y_2 - y_1}{x_2 - x_1} = \frac{4 - 1}{-3 - 5} = \frac{3}{-8} = -\frac{3}{8}\] Now...lets begin......in point slope form...we have \[\large y - y_0 = m(x - x_1)\]

OpenStudy (johnweldon1993):

lets insert 1 of our points...and the slope we found lets use (5 , 1)...we now have \[\large y - 1 = -\frac{3}{8}(x - 5)\] \[\large y - 1 = -\frac{3}{8}x + \frac{15}{8}\] \[\large y = -\frac{3}{8}x + \frac{23}{8}\] would be our equation that we found using point slope form...

OpenStudy (johnweldon1993):

And by going through that whole solving process...we have found the same line in slope intercept form \[\large y = -\frac{3}{8}x + \frac{23}{8}\]